OpenStudy (danieke):

Find the slope and the y-intercept of the equation y − 3(x − 1) = 0. slope = -3 and y-intercept = -3 slope = 3 and y-intercept = 3 slope = -3 and y-intercept = 3 slope = 3 and y-intercept = -3

OpenStudy (alexandervonhumboldt2):

y-3(x-1)=0 y=3(x-1) ok so y=a*(x-x1) where a is slope and a*x1 is y-intercept. a*x1=-3. a*3. thus what?